Trash:

Solve the compound inequality 6b < 42 or 4b + 12 > 8

boredfr:

Ok so looking at \[6b<42\] you have to divide. \[\frac{ 6b }{ 6 }\frac{ 42 }{ 6 }=b<7\]

Trash:

Omg ty for explaining it step by step

boredfr:

Ok so for the other part you have to subtract 12 from both sides so \[4b+12-12>8-12\] which would leave you with \[4b>-4\]

boredfr:

Then you would divide both sides by 4 so \[\frac{ 4b }{ 4 }>\frac{ -4 }{ 4 }\] which would equal \[b>-1\]

boredfr:

So your answer would be \[b<7 \] or \[b>-1\]

boredfr:

Do you understand?

Trash:

Yea i really do tysm

boredfr:

Alr (: yw if you have anyother questions feel free to tag me, also since your question is answered please click the button that says "close"

Trash:

Ok ty and i will